# Anytime-valid, Bayes-assisted, Prediction-Powered Inference

### v1

@@ 7 Discussion
S4.2 Tighter convergence rates for asymptotic confidence sequences
@@
(Tighter version of Theorem 3.4 and Theorem 3.6) Let $(\widehat{\mu}_{t})_{t\geq 1}$ be a consistent sequence of estimators of $\mu$. Assume that there exists a sequence of i.i.d. Gaussian random variables $(W_{i})_{i\geq 1}$, with mean $\mu$ and variance $\sigma^{2}$, such that
$\widehat{\mu}_{t}=\frac{1}{t}\sum_{i=1}^{t}W_{i}+o\left(\frac{1}{\sqrt{t\log t}}\right)\text{ a.s. as }t\to\infty.$ (S41)
Let $(\widehat{\sigma}^{2}_{t})_{t\geq 1}$ be a consistent sequence of estimators of $\sigma^{2}$. Then, for any parameter $\rho>0$, the sequence of intervals $(\mathcal{C}^{\mathrm{\mathtt{N}A}}_{\alpha,t}(\widehat{\mu}_{t},\widehat{\sigma}_{t};\rho))_{t\geq 1}$ forms a $(1-\alpha)$–AsympCS with approximation rate $1/\sqrt{t\log{t}}$ for $\mu$. Additionally, if $\pi$ is a continuous and proper prior density on $\mathbb{R}$ and $\eta_{t}$ is the density (7), then the sequence of intervals $(\mathcal{C}^{\mathrm{\mathtt{B}A}}_{\alpha,t}(\widehat{\mu}_{t},\widehat{\sigma}_{t};\pi))_{t\geq 1}$ forms a $(1-\alpha)$–AsympCS with approximation rate $1/\sqrt{t\log{t}}$ for $\mu$.
@@
(Tighter version of Theorem 3.3 and Theorem 3.5) Let $(Y_{t})_{t\geq 1}$ be a sequence of i.i.d. random variables with mean $\mu$ and such that $\mathbb{E}|Y_{1}|^{2+\delta}<\infty$ for some $\delta>0$. For any $t\geq 1$, let $\overline{Y}_{t}$ be the sample mean, and $\widehat{\sigma}_{t}^{2}$ be the sample variance based on the first $t$ observations. For any parameter $\rho>0$ and any proper continuous prior $\pi$, the sequences of intervals $\left(\mathcal{C}_{\alpha,t}^{\mathrm{\mathtt{N}A}}(\overline{Y}_{t},\widehat{\sigma}_{t};\rho)\right)_{t\geq 1}$ and $\left(\mathcal{C}_{\alpha,t}^{\mathrm{\mathtt{B}A}}(\overline{Y}_{t},\widehat{\sigma}_{t};\pi)\right)_{t\geq 1}$ both form $(1-\alpha)$–AsympCS with approximation rate $1/\sqrt{t\log{t}}$ for $\mu$.
Proof.(Theorem S2) The proof follows the same structure as for the proof of Theorem 3.4 and Theorem 3.6. We explain how to obtain the tighter version of Theorem 3.6; the proof for Theorem 3.4 applies similarly. By assumption, we have, almost surely,
@@
Using Theorem S2, the sequence of intervals $\mathcal{C}^{\mathrm{\mathtt{B}A}}_{\alpha,t}(\overline{W}_{t},\widehat{\sigma}_{t};\pi)=\mathcal{C}^{\mathrm{\mathtt{B}A}}_{\alpha,t}(\widehat{\mu}_{t}-\epsilon_{t},\widehat{\sigma}_{t};\pi)=[\widehat{\mu}_{t}-L_{t}^{\ast},\widehat{\mu}_{t}+U_{t}^{\ast}]$, where
$U_{t}^{\ast}$ $=\frac{\sigma}{\sqrt{t}}\sqrt{\log\left(\frac{t}{2\pi\alpha^{2}}\right)-2\log{\eta}_{t}\left(\frac{\overline{W}_{t}-\varepsilon_{t}}{\sigma}\right)}-\varepsilon_{t},\text{\leavevmode\nobreak\ and}$
$L_{t}^{\ast}$ $=\frac{\sigma}{\sqrt{t}}\sqrt{\log\left(\frac{t}{2\pi\alpha^{2}}\right)-2\log{\eta}_{t}\left(\frac{\overline{W}_{t}-\varepsilon_{t}}{\sigma}\right)}+\varepsilon_{t},$
is an exact confidence sequence for $\mu$. Define $\mathcal{C}^{\mathrm{\mathtt{B}A}}_{\alpha,t}(\widehat{\mu}_{t},\widehat{\sigma}_{t};\pi)=[\widehat{\mu}_{t}-L_{t},\widehat{\mu}_{t}+U_{t}]$, where
@@
$\frac{1}{a_{t}}\left(L_{t}-L_{t}^{*}\right)=\frac{1}{a_{t}}\left[\frac{\widehat{\sigma}_{t}}{\sqrt{t}}\sqrt{\log\left(\frac{t}{2\pi\alpha^{2}}\right)-2\log\eta_{t}\left(\frac{\widehat{\mu}_{t}}{\widehat{\sigma}_{t}}\right)}\right.\\ -\left.\frac{\sigma}{\sqrt{t}}\sqrt{\log\left(\frac{t}{2\pi\alpha^{2}}\right)-2\log\eta_{t}\left(\frac{\bar{W}_{t}-\varepsilon_{t}}{\sigma}\right)}\right]+o(1)$
@@
$\frac{\widehat{\sigma}_{t}}{\sqrt{t}}\sqrt{\log\left(\frac{t}{2\pi\alpha^{2}}\right)-2\log\eta_{t}\left(\frac{\widehat{\mu}_{t}}{\widehat{\sigma}_{t}}\right)}-\frac{\sigma}{\sqrt{t}}\sqrt{\log\left(\frac{t}{2\pi\alpha^{2}}\right)-2\log\eta_{t}\left(\frac{\bar{W}_{t}-\varepsilon_{t}}{\sigma}\right)}$
@@
$=\frac{\frac{\widehat{\sigma}_{t}^{2}}{t}\left[\log\left(\frac{t}{2\pi\alpha^{2}}\right)-2\log\eta_{t}\left(\frac{\widehat{\mu}_{t}}{\widehat{\sigma}_{t}}\right)\right]-\frac{\sigma^{2}}{t}\left[\log\left(\frac{t}{2\pi\alpha^{2}}\right)-2\log\eta_{t}\left(\frac{\bar{W}_{t}-\varepsilon_{t}}{\sigma}\right)\right]}{\frac{\widehat{\sigma}_{t}}{\sqrt{t}}\sqrt{\log\left(\frac{t}{2\pi\alpha^{2}}\right)-2\log\eta_{t}\left(\frac{\widehat{\mu}_{t}}{\widehat{\sigma}_{t}}\right)}+\frac{\sigma}{\sqrt{t}}\sqrt{\log\left(\frac{t}{2\pi\alpha^{2}}\right)-2\log\eta_{t}\left(\frac{\bar{W}_{t}-\varepsilon_{t}}{\sigma}\right)}}$
$\sim\sqrt{\frac{t}{\log t}}\frac{2\sigma}{t}\log\frac{\eta_{t}\left(\frac{\bar{W}_{t}-\varepsilon_{t}}{\sigma}\right)}{\eta_{t}\left(\frac{\widehat{\mu}_{t}}{\widehat{\sigma}_{t}}\right)}=o\left(\frac{1}{\sqrt{t\log t}}\right)$
as, by Lemma S1, we have
@@
${\eta}_{t}\left(\frac{\widehat{\mu}_{t}}{\widehat{\sigma}_{t}}\right)\rightarrow\pi\left(\frac{\mu}{\sigma}\right)\text{\leavevmode\nobreak\ and\leavevmode\nobreak\ }{\eta}_{t}\left(\frac{\bar{W}_{t}-\varepsilon_{t}}{\sigma}\right)\rightarrow\pi\left(\frac{\mu}{\sigma}\right)\text{\leavevmode\nobreak\ a.s. as\leavevmode\nobreak\ }t\rightarrow\infty.$
@@
It follows that $\left(\mathcal{C}_{\alpha,t}^{\mathrm{\mathtt{B}A}}\right)$ is a $(1-\alpha)$–AsympCS with approximation rate $1/\sqrt{t\log{t}}$. ∎
Proof.(Theorem S3) In order to apply Theorem S2 in this setting, we need Equation S41 to be satisfied for the i.i.d. sequence $(Y_{t})_{t\geq 1}$. By KMT strong coupling (Theorem S4), there exists a sequence of i.i.d. Gaussian random variables $(W_{i})_{i\geq 1}$ with mean $\mu$ and variance $\sigma^{2}$ such that, a.s.,

### v2

@@ 7 Discussion
The following lemma will be useful in the proof of Proposition˜1.
Lemma S1.
Let $(U_{i},V_{i})$, $i=1,\ldots,n$, be i.i.d. copies of a pair of random variables $(U,V)$. Assume $\mathbb{E}|U|^{2+\delta}$ and $\mathbb{E}|V|^{2+\delta}<\infty$ for some $0<\delta<1$. Let $\lambda^{\star}=\frac{\mathrm{cov}(U,V)}{\mathrm{var}(U)}$ and $\widehat{\lambda}=\frac{\widehat{\mathrm{cov}}((U_{i},V_{i})_{i=1}^{n})}{\widehat{\mathrm{var}}((U_{i})_{i=1}^{n})}$. Then
$|\lambda^{\star}-\widehat{\lambda}|=o(n^{-\frac{\delta}{2+\delta}})\text{ a.s. as }n\to\infty.$
Proof.
Using the mean value theorem, we obtain
[…]
$+\widehat{\mathrm{cov}}((U_{i},V_{i})_{i=1}^{n})K_{1}\left|\frac{1}{n}\sum_{i=1}^{n}(U_{i}^{2}-\mathbb{E}(U^{2}))\right|+\widehat{\mathrm{cov}}((U_{i},V_{i})_{i=1}^{n})K_{1}K_{2}\left|\overline{U}-\mathbb{E}U\right|$
where $K_{1}$ and $K_{2}$ are two constants, independent of $n$. By assumption, $UV$ and $U^{2}$ have finite moments of order $1+\delta/2$ and $U$ and $V$ have finite moments of order $2+\delta$ (and so of order $1+\delta/2$) thus we can apply Theorem˜S7 with $p=1+\delta/2$ to obtain the result. ∎
S2.2 Confidence sequence for i.i.d. Gaussian variables with known variance
An important step in the proof of all our results is the derivation of an exact confidence sequence for i.i.d. Gaussian variables with known variance. The non-assisted confidence sequence is a well-known result that can be found for instance in [7] or in the proof of Theorem 2.2 in [15]:
@@
Let $W_{i}\overset{\mathrm{iid}}{\sim}\mathcal{N}(\mu,\sigma^{2})$. For any parameter $\rho>0$, the sequence of intervals defined as
$\mathcal{C}^{\mathrm{\tt NA}}_{\alpha,t}(\overline{W}_{t},\sigma;\rho):=\left[\overline{W}_{t}\pm\frac{\sigma}{\sqrt{t}}\sqrt{\left(1+\frac{1}{t\rho^{2}}\right)\log\left(\frac{t\rho^{2}+1}{\alpha^{2}}\right)}\right]$ (S30)
is an exact $1-\alpha$ confidence sequence for $\mu_{0}$, that is
$\Pr\left(\mu\in\mathcal{C}^{\mathrm{\tt NA}}_{\alpha,t}(\overline{W}_{t},\sigma;\rho)\text{ for all }t\geq 1\right)\geq 1-\alpha.$
We also establish such a confidence sequence under a general prior on the mean. While Wang and Ramdas [19, Proposition C.1] propose an exact confidence sequence under a Gaussian prior, we extend their result to any continuous and proper prior.
@@
Let $W_{i}\overset{\mathrm{iid}}{\sim}\mathcal{N}(\mu,\sigma^{2})$ and let $\uppi$ be a continuous and proper on $\mu/\sigma$, then
$\mathcal{C}^{\mathrm{\tt BA}}_{\alpha,t}(\overline{W}_{t},\sigma;\uppi)=\left[\overline{W}_{t}\pm\frac{\sigma}{\sqrt{t}}\sqrt{\log\left(\frac{t}{2\pi\alpha^{2}}\right)-2\log\eta_{t}\left(\frac{\overline{W}_{t}}{\sigma}\right)}\right],$
where $\eta_{t}$ is defined in Equation˜8, is an exact $1-\alpha$ confidence sequence for $\mu_{0}$, that is
$\Pr\left(\mu\in\mathcal{C}^{\mathrm{\tt BA}}_{\alpha,t}(\overline{W}_{t},\sigma;\uppi)\text{ for all }t\geq 1\right)\geq 1-\alpha.$
Of particular interest, we may consider the Gaussian prior $\mathcal{N}(\mu_{0},\tau^{2})$, which gives
$\mathcal{C}^{\mathrm{\tt BA}}_{\alpha,t}(\overline{W}_{t},\sigma;\mathcal{N}(\cdot;\mu_{0},\tau^{2}))=\left[\overline{W}_{t}\pm\frac{\sigma}{\sqrt{t}}\sqrt{\log\left(\frac{t\tau^{2}+1}{\alpha^{2}}\right)+\frac{\left(\overline{W}_{t}/\sigma-\mu_{0}\right)^{2}}{\left(\tau^{2}+1/t\right)}}\right].$
Proof.
The main idea is to apply the methods of mixture with the prior $\uppi$. For each $t\geq 1$, let $p_{t,\mu}(w_{1:t})$ be the joint density of the $W_{1},\ldots,W_{t}$ with respect to $\lambda^{\otimes t}$ where $\lambda$ is the Lebesgue measure, for some unknown mean parameter $\mu\in\mathbb{R}$, where $w_{1:t}=(w_{1},\ldots,w_{t})\in\mathbb{R}^{t}$. To simplify notations, we drop the subscript $t$ and simply write $p_{\mu}(w_{1:t})$ to denote this joint density. We have
$p_{\mu}(w_{1},\ldots,w_{t})=C(\sigma,w_{1},\ldots,w_{t})\times\phi_{t}\left(\frac{\overline{w}_{t}-\mu}{\sigma}\right)$
where $\overline{w}_{t}=\frac{1}{t}\sum_{i=1}^{t}w_{i}$, $C(\sigma,w_{1},\ldots,w_{t})$ does not depend on $\mu$ and $\phi_{t}(w)$ denotes the pdf of a zero-mean Gaussian random variable with variance $1/t$.
For any $\mu\in\mathbb{R},w_{1:t}\in\mathbb{R}^{n}$, let
$\tilde{M}_{t}(\overline{w}_{t},\mu)$ $=\int_{\mathbb{R}}\frac{p_{\mu^{\prime}}(w_{1:t})}{p_{\mu}(w_{1:t})}\uppi\left(\frac{\mu^{\prime}}{\sigma}\right)\frac{d\mu^{\prime}}{\sigma}$ (S31)
$=\int_{\mathbb{R}}\frac{\phi_{t}\left(\frac{\overline{w}_{t}-\mu^{\prime}}{\sigma}\right)}{\phi_{t}\left(\frac{\overline{w}_{t}-\mu}{\sigma}\right)}\uppi\left(\frac{\mu^{\prime}}{\sigma}\right)\frac{d\mu^{\prime}}{\sigma}$ (S32)
$=\frac{\eta_{t}(\overline{w}_{t}/\sigma)}{\phi_{t}((\overline{w}_{t}-\mu)/\sigma)}.$ (S33)
We define $M_{t}(\mu^{\prime})=\frac{p_{\mu^{\prime}}(W_{1:t})}{p_{\mu}(W_{1:t})}=\exp\left(-\frac{1}{2\sigma^{2}}\left(\sum_{i=1}^{t}(W_{i}-\mu^{\prime})^{2}-\sum_{i=1}^{t}(W_{i}-\mu)^{2}\right)\right).$ We consider $(\mathcal{F}_{t})_{t\in\mathbb{N}}$, the filtration adapted to the sequence of random variable $(W_{t})_{t\in\mathbb{N}}$. For every $\mu^{\prime}\in\mathbb{R}$ we have
$M_{t}(\mu^{\prime})$ $=M_{t-1}(\mu^{\prime})\times\exp\left(\frac{1}{2\sigma^{2}}(\mu^{2}-{\mu^{\prime}}^{2})\right)\exp\left(-\frac{W_{t}}{\sigma^{2}}(\mu-\mu^{\prime})\right),$
and so
$\mathbb{E}\left[M_{t}(\mu^{\prime})\mid\mathcal{F}_{t-1}\right]$ $=M_{t-1}(\mu^{\prime})\times\exp\left(\frac{1}{2\sigma^{2}}(\mu^{2}-{\mu^{\prime}}^{2})\right)\mathbb{E}\exp\left(-\frac{W_{t}}{\sigma^{2}}(\mu-\mu^{\prime})\right)$
$=M_{t-1}(\mu^{\prime})\times\exp\left(\frac{1}{2\sigma^{2}}(\mu^{2}-{\mu^{\prime}}^{2})\right)\exp\left(-\mu\frac{\mu-\mu^{\prime}}{\sigma^{2}}+\frac{1}{2\sigma^{2}}(\mu-{\mu^{\prime}})^{2}\right)$
$=M_{t-1}(\mu^{\prime}).$
Hence, $\{(M_{t}(\mu^{\prime}))_{t\in\mathbb{N}},\mu^{\prime}\in\mathbb{R}\}$ is a family of martingale with respect to the adapted filtration $(\mathcal{F}_{t})_{t\in\mathbb{N}}$.
$M_{t}(\mu^{\prime})$ is clearly continuous in $W_{i}$ for all $i\in\{1,...,t\}$ and for $\mu^{\prime}\in\mathbb{R}$. Hence, it is $\mathcal{F}_{t}\otimes\mathcal{B}(\mathbb{R})$-measurable. Similarly, $\mathbb{E}\left[M_{t}(\mu^{\prime})\mid\mathcal{F}_{t-1}\right]=M_{t-1}(\mu^{\prime})$ is $\mathcal{F}_{t-1}\otimes\mathcal{B}(\mathbb{R})$-measurable.
Finally, we have $M_{t}(\mu^{\prime})\geq 0$ $\Pr\otimes\lambda$-almost everywhere, where $\lambda$ is the Lebesgue measure on $\mathbb{R}$ (or any other measure dominated by the Lebesgue measure on $\mathbb{R}$). Then, by Proposition˜S8, $(\tilde{M}_{t}(\overline{W}_{t},\mu))_{t\geq 1}$ is a nonnegative martingale with respect to the adapted filtration $(\mathcal{F}_{t})_{t\in\mathbb{N}}$. So by Ville’s inequality, we have :
$\Pr\left(\tilde{M}_{t}(\overline{W}_{t},\mu)\leq\frac{1}{\alpha}\text{ for all }t\geq 1\right)\geq 1-\alpha.$
It follows that the sequence
$\mathcal{C}^{\mathrm{\tt BA}}_{\alpha,t}(\overline{W}_{t},\sigma;\uppi)=\left\{\mu~|~\tilde{M}_{t}(\overline{W}_{t},\mu)\leq\frac{1}{\alpha}\right\}$
is an exact $1-\alpha$ confidence sequence for $\mu$. Finally,
$\tilde{M}_{t}(\overline{W}_{t},\mu)\leq\frac{1}{\alpha}$ $\Longleftrightarrow\exp\left(\frac{t}{2\sigma^{2}}(\mu-\overline{W}_{t})^{2}\right)\eta_{t}\left(\frac{\overline{W}_{t}}{\sigma}\right)\frac{\sqrt{2\pi}}{\sqrt{t}}\leq\frac{1}{\alpha}$
$\Longleftrightarrow\frac{t}{2\sigma^{2}}(\mu-\overline{W}_{t})^{2}+\log\left({\eta}_{t}\left(\frac{\overline{W}_{t}}{\sigma}\right)\frac{\sqrt{2\pi}}{\sqrt{t}}\right)\leq-\log(\alpha)$
$\Longleftrightarrow(\mu-\overline{W}_{t})^{2}\leq\frac{\sigma^{2}}{t}\left(-2\log\left({\eta}_{t}\left(\frac{\overline{W_{t}}}{\sigma}\right)\right)-\log\left(\frac{2\pi\alpha^{2}}{t}\right)\right).$
S2.3 Optimal control-variate parameter for PPI++
The next proposition follows from an application of Proposition˜1 to the PPI++ estimators, identifying the value of the optimal control variate parameter in this case.
Proposition S10 (Asymptotics for PPI++).
$\lambda_{\theta}^{\star}=\frac{\mathrm{cov}(\ell_{\theta}^{\prime}(X_{1},Y_{1}),\ell_{\theta}^{\prime}(X_{1},f(X_{1})))}{\mathrm{var}(\ell_{\theta}^{\prime}(X_{1},f(X_{1})))}.$ (S34)
Then, for any $\theta$, almost surely as $n\to\infty$,
$\widehat{\Delta}^{\mathrm{\texttt{PP+}}}_{\theta,n}$ $=\frac{1}{n}\sum_{i=1}^{n}\left(\ell_{\theta}^{\prime}(X_{i},Y_{i})-\ell_{\theta}^{\prime}(X_{i},f(X_{i}))\right)-(\lambda^{\star}_{\theta}-1)\left(\frac{1}{n}\left[\sum_{i=1}^{n}\ell_{\theta}^{\prime}(X_{i},f(X_{i}))\right]-\widehat{m}_{\theta}\right)$
$~~~~+o\left(\sqrt{\frac{\log\log n}{n}}\right).$ (S36)
Additionally, in the case of the squared loss,
$\widehat{\theta}_{n}^{\mathrm{\texttt{PP+}}}$ $=\frac{1}{n}\sum_{i=1}^{n}Y_{i}-\lambda^{\star}_{0}\left(\frac{1}{n}\sum_{i=1}^{n}f(X_{i})-\frac{1}{N_{n}}\sum_{j=1}^{N_{n}}f(\widetilde{X}_{j})\right)+o\left(\sqrt{\frac{\log\log n}{n}}\right),$ (S37)
with $\lambda^{\star}_{0}=\mathrm{cov}(Y_{1},f(X_{1}))/\mathrm{var}(f(X_{1}))$.
Appendix S3 Proofs
S3.1 Proofs of Theorem˜1 and Theorem˜2
Theorem˜1 is a corollary of Theorem˜2, so we start by proving Theorem˜2.
S3.1.1 Proof of Theorem˜2
By assumption, we have, almost surely,
@@
Using Theorem˜S10, the sequence of intervals $\mathcal{C}^{\mathrm{\tt NA}}_{\alpha,t}(\overline{W}_{t},\sigma;\rho)=\mathcal{C}^{\mathrm{\tt NA}}_{\alpha,t}(\widehat{\mu}_{t}-\varepsilon_{t},\sigma;\rho)=[\widehat{\mu}_{t}-L_{t}^{\ast},\widehat{\mu}_{t}+U_{t}^{\ast}]$, where
$U_{t}^{\ast}$ $=\frac{\sigma}{\sqrt{t}}\sqrt{\left(1+\frac{1}{t\rho^{2}}\right)\log\left(\frac{t\rho^{2}+1}{\alpha^{2}}\right)}-\varepsilon_{t},\text{~and}$
$L_{t}^{\ast}$ $=\frac{\sigma}{\sqrt{t}}\sqrt{\left(1+\frac{1}{t\rho^{2}}\right)\log\left(\frac{t\rho^{2}+1}{\alpha^{2}}\right)}+\varepsilon_{t},$
is an exact confidence sequence for $\mu$. We have $\mathcal{C}^{\mathrm{\tt NA}}_{\alpha,t}(\widehat{\mu}_{t},\widehat{\sigma}_{t};\rho)=[\widehat{\mu}_{t}-L_{t},\widehat{\mu}_{t}+U_{t}]$, where
$U_{t}=L_{t}=\frac{\widehat{\sigma}_{t}}{\sqrt{t}}\sqrt{\left(1+\frac{1}{t\rho^{2}}\right)\log\left(\frac{t\rho^{2}+1}{\alpha^{2}}\right)}.$
Let $a_{t}=1/\sqrt{t\log t}$. Then,
$\frac{1}{a_{t}}\left(L_{t}-L_{t}^{\star}\right)=\frac{1}{a_{t}}\left[\frac{\widehat{\sigma}_{t}}{\sqrt{t}}\sqrt{\left(1+\frac{1}{t\rho^{2}}\right)\log\left(\frac{t\rho^{2}+1}{\alpha^{2}}\right)}\right.\\ -\left.\frac{\sigma}{\sqrt{t}}\sqrt{\left(1+\frac{1}{t\rho^{2}}\right)\log\left(\frac{t\rho^{2}+1}{\alpha^{2}}\right)}\right]+o(1).$
We have
$\frac{\widehat{\sigma}_{t}}{\sqrt{t}}\sqrt{\left(1+\frac{1}{t\rho^{2}}\right)\log\left(\frac{t\rho^{2}+1}{\alpha^{2}}\right)}-\frac{\sigma}{\sqrt{t}}\sqrt{\left(1+\frac{1}{t\rho^{2}}\right)\log\left(\frac{t\rho^{2}+1}{\alpha^{2}}\right)}$
$\sim(\widehat{\sigma}_{t}-\sigma)\times\sqrt{\frac{\log t}{t}}=o\left(\frac{1}{\sqrt{t\log t}}\right).$
Hence $\frac{1}{a_{t}}\left(L_{t}-L_{t}^{\star}\right)=o(1)$. Similarly, $\frac{1}{a_{t}}\left(U_{t}-U_{t}^{\star}\right)=o(1)$. It follows that $\left(\mathcal{C}_{\alpha,t}^{\mathrm{\tt NA}}\right)$ is a $(1-\alpha)$–AsympCS with approximation rate $1/\sqrt{t\log{t}}$.
S3.1.2 Proof of Theorem˜1
In order to apply Theorem˜2 in this setting, we need Equation˜7 to be satisfied for the i.i.d. sequence $(Y_{t})_{t\geq 1}$. By KMT strong coupling (Theorem˜S8), there exists a sequence of i.i.d. Gaussian random variables $(W_{i})_{i\geq 1}$ with mean $\mu$ and variance $\sigma^{2}$ such that, a.s.,
$\overline{Y}_{t}=\frac{1}{t}\sum_{i=1}^{t}W_{i}+\varepsilon_{t}\text{~~~where~~~}\varepsilon_{t}=o\left(\frac{1}{t^{1-1/(2+\delta)}}\right)=o\left(\frac{1}{\sqrt{t\log t}}\right).$
We also need to satisfied the condition on the variance. Under $\mathbb{E}[|Y|^{2+\delta}]<\infty$, the Marcinkiewicz-Zygmund strong law of large numbers (Theorem˜S7) with $p=1+\delta/2\in(1,2)$ yields a polynomial a.s. rate for $\overline{Y}_{t}$ and $\overline{Y^{2}_{t}}$; consequently $|\widehat{\sigma}_{t}-\sigma|=o(t^{-\gamma})$ for some $\gamma>0$, which implies
$|\widehat{\sigma}_{t}-\sigma|=o\left(\frac{1}{\log t}\right)\text{ a.s. as }t\to\infty.$
The result follows.
S3.2 Proofs of Theorem˜3 and Theorem˜4
The following lemma is a direct consequence of Theorem 8.14(b) p. 242 [20].
Lemma S2.
Let $\uppi$ be a proper and continuous probability density function on $\mathbb{R}^{d}$. Let $\left(Z_{t}\right)_{t\geq 1}$ be a sequence of random vectors in $\mathbb{R}^{d}$, with $Z_{t}\rightarrow c$ a.s. as $t\rightarrow\infty$. Let
$\eta_{t}(z)=\int_{\mathbb{R}^{d}}\mathcal{N}(z;\zeta,I_{d}/t)\uppi(\zeta)d\zeta.$
Then
${\eta}_{t}(Z_{t})\rightarrow\uppi(c)\text{ almost surely as }t\rightarrow\infty.$
Theorem˜3 is a corollary of Theorem˜4, so we start by proving Theorem˜4.
S3.2.1 Proof of Theorem˜4
By assumption, we have, almost surely,
$\widehat{\mu}_{t}=\overline{W}_{t}+\varepsilon_{t},$
Using Theorem˜S11, the sequence of intervals $\mathcal{C}^{\mathrm{\tt BA}}_{\alpha,t}(\overline{W}_{t},\sigma_{;}\uppi)=\mathcal{C}^{\mathrm{\tt BA}}_{\alpha,t}(\widehat{\mu}_{t}-\varepsilon_{t},\sigma;\uppi)=[\widehat{\mu}_{t}-L_{t}^{\ast},\widehat{\mu}_{t}+U_{t}^{\ast}]$, where
$U_{t}^{\ast}$ $=\frac{\sigma}{\sqrt{t}}\sqrt{\log\left(\frac{t}{2\pi\alpha^{2}}\right)-2\log{\eta}_{t}\left(\frac{\widehat{\mu}_{t}-\varepsilon_{t}}{\sigma}\right)}-\varepsilon_{t},\text{~and}$
$L_{t}^{\ast}$ $=\frac{\sigma}{\sqrt{t}}\sqrt{\log\left(\frac{t}{2\pi\alpha^{2}}\right)-2\log{\eta}_{t}\left(\frac{\widehat{\mu}_{t}-\varepsilon_{t}}{\sigma}\right)}+\varepsilon_{t},$
is an exact confidence sequence for $\mu$. We have $\mathcal{C}^{\mathrm{\tt BA}}_{\alpha,t}(\widehat{\mu}_{t},\widehat{\sigma}_{t};\uppi)=[\widehat{\mu}_{t}-L_{t},\widehat{\mu}_{t}+U_{t}]$, where
@@
$\frac{1}{a_{t}}\left(L_{t}-L_{t}^{\star}\right)=\frac{1}{a_{t}}\left[\frac{\widehat{\sigma}_{t}}{\sqrt{t}}\sqrt{\log\left(\frac{t}{2\pi\alpha^{2}}\right)-2\log\eta_{t}\left(\frac{\widehat{\mu}_{t}}{\widehat{\sigma}_{t}}\right)}\right.\\ -\left.\frac{\sigma}{\sqrt{t}}\sqrt{\log\left(\frac{t}{2\pi\alpha^{2}}\right)-2\log\eta_{t}\left(\frac{\widehat{\mu}_{t}-\varepsilon_{t}}{\sigma}\right)}\right]+o(1)$
@@
$\frac{\widehat{\sigma}_{t}}{\sqrt{t}}\sqrt{\log\left(\frac{t}{2\pi\alpha^{2}}\right)-2\log\eta_{t}\left(\frac{\widehat{\mu}_{t}}{\widehat{\sigma}_{t}}\right)}-\frac{\sigma}{\sqrt{t}}\sqrt{\log\left(\frac{t}{2\pi\alpha^{2}}\right)-2\log\eta_{t}\left(\frac{\widehat{\mu}_{t}-\varepsilon_{t}}{\sigma}\right)}$
@@
$=\frac{\frac{\widehat{\sigma}_{t}^{2}}{t}\left[\log\left(\frac{t}{2\pi\alpha^{2}}\right)-2\log\eta_{t}\left(\frac{\widehat{\mu}_{t}}{\widehat{\sigma}_{t}}\right)\right]-\frac{\sigma^{2}}{t}\left[\log\left(\frac{t}{2\pi\alpha^{2}}\right)-2\log\eta_{t}\left(\frac{\widehat{\mu}_{t}-\varepsilon_{t}}{\sigma}\right)\right]}{\frac{\widehat{\sigma}_{t}}{\sqrt{t}}\sqrt{\log\left(\frac{t}{2\pi\alpha^{2}}\right)-2\log\eta_{t}\left(\frac{\widehat{\mu}_{t}}{\widehat{\sigma}_{t}}\right)}+\frac{\sigma}{\sqrt{t}}\sqrt{\log\left(\frac{t}{2\pi\alpha^{2}}\right)-2\log\eta_{t}\left(\frac{\widehat{\mu}_{t}-\varepsilon_{t}}{\sigma}\right)}}$
$\sim(\widehat{\sigma}_{t}-\sigma)\times\sqrt{\frac{\log t}{t}}=o\left(\frac{1}{\sqrt{t\log t}}\right)$
as, by Lemma˜S2, we have
@@
${\eta}_{t}\left(\frac{\widehat{\mu}_{t}}{\widehat{\sigma}_{t}}\right)\rightarrow\uppi\left(\frac{\mu}{\sigma}\right)\text{~and~}{\eta}_{t}\left(\frac{\widehat{\mu}_{t}-\varepsilon_{t}}{\sigma}\right)\rightarrow\uppi\left(\frac{\mu}{\sigma}\right)\text{~a.s. as~}t\rightarrow\infty.$
@@
It follows that $\left(\mathcal{C}_{\alpha,t}^{\mathrm{\tt BA}}\right)$ is a $(1-\alpha)$–AsympCS with approximation rate $1/\sqrt{t\log{t}}$.
S3.2.2 Proof of Theorem˜3
The result follows from Theorem˜4 by applying the same reasoning as for the proof of Theorem˜1.
S3.3 Proof of Theorem˜5
The idea of the proof is as follows. Recall that the intervals $(\mathcal{C}_{\alpha,t})$ of interest are approximations of an exact CS $(\mathcal{C}^{\star}_{\alpha,t})$. For any $\alpha^{\prime}>\alpha$, the narrower exact CS $(\mathcal{C}^{\star}_{\alpha^{\prime},t})$ is eventually (a.s.) contained in $(\mathcal{C}_{\alpha,t})$ for all large $t$. A standard sandwiching argument using this eventual containment yields the desired asymptotic Type-I control.
Let $a_{t}=1/\sqrt{t\log t}$. For each construction (non-assisted/Bayes-assisted), the sequence of intervals of interest are of the form $\mathcal{C}_{\alpha,t}=[\widehat{\mu}_{t}\pm U_{\alpha,t}]$ where
$U_{\alpha,t}=\frac{\widehat{\sigma}_{t}}{\sqrt{t}}\sqrt{\left(1+\frac{1}{t\rho^{2}}\right)\log\left(\frac{t\rho^{2}+1}{\alpha^{2}}\right)}$
for the non-assisted case, and
$U_{\alpha,t}=\frac{\widehat{\sigma}_{t}}{\sqrt{t}}\sqrt{\log\left(\frac{t}{2\pi\alpha^{2}}\right)-2\log{\eta}_{t}\left(\frac{\widehat{\mu}_{t}}{\widehat{\sigma}_{t}}\right)}$
for the Bayes-assisted case. $\mathcal{C}_{\alpha,t}$ approximates a reference exact CS of the form $\mathcal{C}^{\star}_{\alpha,t}=[\widehat{\mu}_{t}-L^{\star}_{\alpha,t},\widehat{\mu}_{t}+U^{\star}_{\alpha,t}]$, where